OpenStudy (anonymous):

Help If f(x)=-3x+3, then f^-1(x)=

OpenStudy (welshfella):

first find x in terms of f(x) can you do that?

OpenStudy (anonymous):

no

OpenStudy (welshfella):

ok its like solving for x f(x) = -3x + 3 add 3x to both sides f(x) + 3x = -3x + 3x + 3 the tems in x disappear on right side 3x + f(x) = 3 now subtract f(x) from both sides 3x = 3 - f(x) finally divide both sides by 3:- x = (3 - f(x)) / 3 replace f(x) by x and we have the inverse of f(x) f-1 (x) = (3 - x)/3

OpenStudy (welshfella):

that is it

OpenStudy (anonymous):

so solve for the bottom one or is that the answer

OpenStudy (welshfella):

That is the answer

OpenStudy (welshfella):

f^-1 (x) is the inverse of f(x)

OpenStudy (anonymous):

Thank you

OpenStudy (welshfella):

yw

OpenStudy (anonymous):

but how would you do this one If f(x)=3x-7, then f^-1(x)=

OpenStudy (welshfella):

in a similar way to the first

OpenStudy (anonymous):

i tried but i don't get it